# Mixture-of-Recursions: Learning Dynamic Recursive Depths for Adaptive Token-Level Computation

### v2

@@ 7.1 Limitations and Future Works
Recent studies have highlighted the redundancy within reasoning chains and address it by applying token-level adaptive computation, like early-exit mechanisms (Yang et al., 2025; Jiang et al., 2025; Dai et al., 2025). Our MoR framework inherently enables latent reasoning by adaptively determining the necessary recursion depth for individual tokens. Therefore, a crucial future work involves exploring how the router can dynamically learn to adjust to the necessity of chain-of-thought (CoT) chains when post-trained on actual reasoning datasets. Developing advanced routing strategies that explicitly align recursion depth with reasoning complexity may enhance reasoning accuracy, computational efficiency, and even interpretability.
@@
Our experiments have been limited to models with up to 1.7 billion parameters due to compute constraints. The natural next step is to train MoR models at larger scales (over 3 billion parameters) on substantially larger corpora. To reduce overall pre-training costs, we could also explore continued pre-training (i.e., uptraining), starting from existing pre-trained vanilla LLM checkpoints. As future work, we plan to investigate MoR performance using various initialization strategies for recursive models, as explored in prior work (Bae et al., 2024). Additionally, to ensure a fair scalability comparison, we need to account for potential performance degradation in Recursive Transformers during post-training for early-exiting (Bae et al., 2024) and incorporate inference throughput constraints for Vanilla Transformers.
@@
Given MoR’s token-level adaptive recursion, we can further optimize computation by integrating structured sparsity. This approach allows for the selective activation of subnetworks or parameters (Liu et al., 2023b), dynamically pruning unnecessary computations at both the token and layer levels (Raposo et al., 2024; Elhoushi et al., 2024). This investigation into sparse model designs promises significant efficiency improvements. We believe many sparsity-based techniques, such as pruning (Han et al., 2015) or quantization (Jacob et al., 2018), are highly complementary to MoR. This will provide deeper insights into effective sparse architectures within recursive models, offering promising directions for future research.
Expansion to multimodal and non-text domains.
@@ 7.2 Acknowledgements
We thank Jacob Eisenstein for valuable feedback on an earlier version of the paper. We also thank Seungyeon Kim, Mostafa Elhoushi, Sangdoo Yun for helpful conversations. Finally, we thank the Google Cloud Platform for awarding Google Cloud credits for this project.
@@ Appendix B Experimental Setup
Model Architecture Details
@@
Appendix C Expanded Results of IsoFLOP Analysis
In the main paper (§3.2), we compared Vanilla, Recursive and our Mixture-of-Recursions (MoR) models under matched training compute. Four base model capacities were studied—135M, 360M, 730M and 1.7B parameters. For recursive and MoR models, we fix the recursion count to $N_{r}\!=\!3$, so the number of unique parameters is roughly one-third of the vanilla counterpart. Each architecture is trained once for the largest compute budget (16.5EB)6661EB = $\,{=}\,10^{18}$ floating-point operations. and the resulting checkpoint is re-used to obtain the 5EB and 2EB variants, as detailed below.
@@ Appendix C Expanded Results of IsoFLOP Analysis
Trapezoid learning-rate schedule with checkpoint reuse.
@@
$\eta(t)=\begin{cases}\frac{t}{w}\,\eta_{\max},&0\leq t<w\quad\text{(warm\mbox{-}up)},\\[4.0pt] \eta_{\max},&w\leq t<p\quad\text{(plateau)},\\[2.0pt] \eta_{\max}\!\Bigl{(}1-\tfrac{t-p}{d}\Bigr{)},&p\leq t<p+d\quad\text{(cool\mbox{-}down)},\end{cases}$
where $w$ denotes the warm-up interval, $p-w$ is the constant-LR plateau, and $d$ represents the cool-down segment. This stable phase allows us to efficiently manage experiments by saving intermediate checkpoints and then running additional cool-down steps from those points, according to each budget. For the warm-up we allocate $5\,\%$ of the total training steps of the smallest budget (2EB), and we set the cool-down steps to $20\,\%$ of the total training steps for each corresponding budget.
@@
Table 7 reports NLL on FineWeb-Edu validation set and few-shot accuracy on six benchmarks. Our findings reveal a clear trend: more compute consistently leads to better models, evidenced by lower NLL and improved accuracy with higher FLOPs. However, weight-sharing alone in recursive models degraded performance compared to Vanilla, a clear trade-off for the reduced parameters. Crucially, token-routed MoR models overcome this shortcoming, catching up to and then surpassing Vanilla models from 360M parameters upward, all while utilizing only one-third of the parameters. This performance advantage persists at 730M and 1.7B parameter scales.
Table 7: Detailed results of isoFLOP analysis across three compute budgets. We evaluate negative log-likelihood (NLL) on the FineWeb-Edu validation set and few-shot accuracy on six downstream tasks for four base model sizes (135M, 360M, 730M, 1.7B). Each model was initially trained up to 16.5EB and sliced back to 5EB and 2EB via mid-training checkpoints using a trapezoid learning-rate schedule. All models used three recursion steps. For MoR models, we use expert-choice routing and recursion-wise caching mechanisms. We highlight the best-performing model in each setting in gray.
@@
Appendix D Details of Experimental Settings for Throughput Measurement
@@
Appendix E Expanded Results of Parameter Sharing Strategy
This section complements the ablation in §4.1 by providing the full quantitative panorama behind 4(b). We revisit the four weight-tying schemes—Cycle, Sequence, Middle-Cycle, and Middle-Sequence—on two base model scales (135M and 360M non-embedding parameters) and two different recursion depths ($N_{r}=2$ and 3). All models were trained from scratch for 10B tokens under identical optimization hyperparameters. Validation NLL on FineWeb-Edu and averaged few-shot accuracy over six benchmarks are summarized in Table 8.
Middle-Cycle is consistently the safest choice.
@@ Appendix E Expanded Results of Parameter Sharing Strategy
Figure 6: Validation negative log-likelihood (lower is better) on FineWeb-Edu for four parameter-sharing strategies. Bars are grouped by model capacity (135M vs. 360M parameters). Middle-Cycle consistently attains the lowest NLL, with its margin widening as either model size or depth increases. The horizontal dashed lines mark the untied (non-sharing) baselines: the lower red line represents the full capacity model, while the upper black line represents a parameter-matched reduced model with a footprint equal to the unique trainable parameter sizes of the recursive model.
Behavior under continued pre-training (up-training).
Table 9 extends the study by “up-training” models—continuing from open-sourced SmolLM (Allal et al., 2024) checkpoints for an additional 5B tokens. Both Middle strategies demonstrate superior performance across all settings, and notably, they significantly outperform the reduced baseline models that are initialized in the same manner but without recursion. The other strategies reach a performance plateau earlier, suggesting that they have limited room for further improvement in capacity.
@@
Appendix F Expanded Results of Design Choices for Router
F.1 Details of Design Configurations
@@
F.2 Router Performance Evaluation Metrics
@@
$\mathrm{MaxVio}=\frac{\max_{i}\mathrm{Load}_{i}-\overline{\mathrm{Load}}_{i}}{\overline{\mathrm{Load}}_{i}},$
@@
F.3 Extended Evaluation Results of Router Designs
@@
Appendix G Expanded Results of KV Cache Sharing Mechanism
G.1 Key Value Representation Trends in Recursive Transformers
Sharing KV caches across model depths has emerged as a promising approach to improve inference throughput in Vanilla Transformers (Brandon et al., 2024). This technique can reduce the memory footprint required for KV caches, enabling larger inference batch sizes. Significant speedups can be achieved by skipping the KV projection and even prefill operations at shared depths in specific designs (Sun et al., 2024). Due to the high degree of freedom in Vanilla models—where trainable parameters can be well optimized for shared caches—these models exhibit only marginal performance drops when KV caches are shared between adjacent layers. In contrast, Recursive Transformers have far fewer parameters available for being optimized to tied key-value states. Nevertheless, we hypothesize that similar patterns may emerge between shared attention blocks. To investigate this, we decomposed the KV states from pretrained Recursive Transformers into magnitude and directional components.
@@ G.1 Key Value Representation Trends in Recursive Transformers
G.2 Performance Comparison of KV Sharing Strategy
@@
Table 12: Comparison of KV cache sharing strategies across Vanilla, Recursive, and MoR Transformers. Models are pretrained on 10B tokens of FineWeb-Edu, and evaluated using negative log-likelihood (NLL) on train set and few-shot accuracy across benchmarks. KV sharing denotes use of recursive KV sharing mechanism. If MoR is mentioned without further specification of KV sharing strategy, it implies the use of the recursion-wise caching strategy. $\text{}^{\dagger}$t indicates training with hybrid KV sharing that leverages shared caches for inactive positions while updating active ones through actual computation.
@@
Appendix H Expanded Qualitative Results
H.1 Analysis on Adaptive Computation Paths
@@
Table 14: Visualization of the recursion depth for each subword token, with colors representing the number of recursion steps: 1, 2, and 3. Each row corresponds to a single sample, offering a clear illustration of the token-level recursion distribution in practice. We use an MoR model with $N_{r}=3$, auxiliary loss, and recursion-wise KV caching. This model is built on a 360M parameter base and trained on 30B tokens.
@@
H.2 Analysis on Router Weights

### v3

@@ 7.1 Limitations and Future Works
Recent studies have highlighted the redundancy within reasoning chains and address it by applying token-level adaptive computation, like early-exit mechanisms (Yang et al., 2025; Jiang et al., 2025; Dai et al., 2025). Our MoR framework inherently enables latent reasoning by adaptively determining the necessary recursion depth for individual tokens. Therefore, a crucial future work involves exploring how the router can dynamically learn to adjust to the necessity of chain-of-thought (CoT) chains when post-trained on actual reasoning datasets. Developing advanced routing strategies that explicitly align recursion depth with reasoning complexity may enhance reasoning accuracy, computational efficiency, and even interpretability for deliberative reasoning process.
@@
Due to current compute constraints, our experiments have been limited to models of up to 1.7 billion parameters. The next step is to train Mixture-of-Recursions (MoR) models at a larger scale (over 3 billion parameters) on substantially larger corpora. To enhance the scalability of the MoR architecture, we can first increase the size of the non-shared block. A potential follow-up work could introduce depth-specific LoRA (Bae et al., 2024) or experts (i.e., incorporating Mixture-of-Experts principles) and leverage expert parallelism (Rajbhandari et al., 2022) to efficiently compute the input at each depth in parallel, which is expected to improve model quality without huge speed bottlenecks. Furthermore, to reduce overall pre-training costs, we could also explore continued pre-training (uptraining), starting from existing pre-trained vanilla LLM checkpoints. As future work, we plan to investigate MoR performance using various initialization strategies for recursive models, as explored in prior work (Bae et al., 2024).
@@
Given MoR’s token-level adaptive recursion, we can further optimize computation by integrating structured sparsity. This approach allows for the selective activation of subnetworks or parameters (Liu et al., 2023b), dynamically pruning unnecessary computations at both the token and layer levels (Raposo et al., 2024; Elhoushi et al., 2024). This investigation into sparse model designs promises significant efficiency improvements. We believe many sparsity-based techniques, such as pruning (Han et al., 2015) or quantization (Jacob et al., 2018), are highly complementary to our MoR framework. This will provide deeper insights into effective sparse architectures within recursive models, offering promising directions for future research.
Expansion to multimodal and non‑text domains.
@@ 7.2 Acknowledgements
We thank Jacob Eisenstein for valuable feedback on an earlier version of the paper. We also acknowledge the support and helpful conversations from Seungyeon Kim, Mostafa Elhoushi, Pascal Vincent, Irina Rish, Sangdoo Yun, and Baeseong Park. Finally, we thank the Google Cloud Platform for awarding Google Cloud credits for this project, and Google’s research grant project for its support.
This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) ([RS-2019-II190075, Artificial Intelligence Graduate School Program (KAIST), 5%], [No. RS-2024-00457882, AI Research Hub Project, 5%], and [No. 2022-0-00871, Development of AI Autonomy and Knowledge Enhancement for AI Agent Collaboration, 90%]). This research was also enabled in part by computational resources, software, and technical assistance provided by Mila, ULaval, Calcul Québec, and Digital Research Alliance of Canada.
@@ Appendix B Experimental Setup
Model architecture details.
@@
Appendix C Expanded Results of Main Experiments
C.1 Performance Validation at 1.7B Scale
We validate our models by scaling up the size to 1.7B, all under the same FLOPs budget. As shown in Table 7, the MoR model with expert-choice routing shows strong performance, but the vanilla model performs slightly better. While this could be due to the differences in the optimal number of tokens associated with each model’s optimal scaling policy, it might also signal that the current architecture design of MoR is not suitable for scaling. Nevertheless, we may be able to further improve performance by increasing the number of non-shared blocks (despite reduced efficiency), applying depth-specific LoRA or MoE to each recursion (Bae et al., 2024), or uptraining with initialization from a well-pretrained checkpoint. Further validation with larger sizes or training on more tokens is required.
[…]
MoR Recursion Pretrain Few-shot Accuracy $\uparrow$
Models Type KV Share $N_{r}$ Param FLOPs $N_{tok}$ LD HS PQ WG ARC MMLU Avg
Vanilla - - - - 1.61B 68.5 20B 40.8 49.4 70.6 54.8 47.4 30.2 48.9
$\text{Recursive}^{\dagger}$ - - M-Cyc 2 0.87B 68.5 18B 37.3 46.5 68.9 52.6 44.2 29.6 46.5
- - M-Cyc 3 0.67B 68.5 20B 36.4 45.3 69.5 52.7 43.9 29.1 46.2
Expert Cache M-Cyc 2 0.87B 68.5 26B 41.1 47.5 70.0 55.6 46.0 30.3 48.4
Expert Cache M-Cyc 3 0.67B 68.5 27B 37.2 46.6 69.1 53.7 44.1 29.7 46.7
MoR (ours) Token Cache M-Cyc 3 0.67B 68.5 30B 35.6 43.2 68.1 53.0 43.4 29.0 45.4
C.2 Increasing Recursion Depth under Fixed Parameters
Table 8 illustrates the performance change when the number of recursions ($N_{r}$) is increased while the number of unique parameters is fixed at 118M. As the number of recursions rises from 1 (Vanilla) to 2 and 3 (MoR), the negative log-likelihood decreases on both the train and validation sets, and the average few-shot accuracy improves. This demonstrates that the MoR model can effectively scale performance by merely increasing the recursion steps, even with a fixed model size.
Table 8: Comparison of MoR with different numbers of recursions under the same number of parameters. Vanilla equals MoR with recursion 1. All models are trained on FineWeb-Edu with 10B tokens, and we apply the Middle-Cycle parameter sharing for MoR models. We report negative log-likelihood (NLL) on the train and validation sets and few-shot accuracy across six benchmarks.
Pretrain Recursion NLL $\downarrow$ Few-shot Accuracy $\uparrow$
Models N-Emb $N_{L}$ $N_{tok}$ $N_{r}$ Share Train Valid LD HS PQ WG ARC MMLU Avg
Vanilla 118M 12 10B 1 - 2.9182 2.9678 25.81 33.08 61.81 51.46 38.10 26.64 39.48
MoR 118M 1+10+1 10B 2 M-Cyc 2.8767 2.9205 27.09 33.84 62.13 53.04 36.77 26.88 39.96
MoR 118M 1+10+1 10B 3 M-Cyc 2.8667 2.9111 27.38 34.60 63.17 51.46 37.15 26.83 40.10
Appendix D Expanded Results of IsoFLOP Analysis
In the main paper (§3.2), we compared Vanilla, Recursive and our Mixture‑of‑Recursions (MoR) models under matched training compute. Four base model capacities were studied—135M, 360M, 730M and 1.7B parameters. For recursive and MoR models, we fix the recursion count to $N_{r}\!=\!3$, so the number of unique parameters is roughly one‑third of the vanilla counterpart. Each architecture is trained once for the largest compute budget (16.5EB)6661EB = $10^{18}$ floating‑point operations. and the resulting checkpoint is re‑used to obtain the 5EB and 2EB variants, as detailed below.
@@ Appendix D Expanded Results of IsoFLOP Analysis
Trapezoid learning‑rate schedule with checkpoint reuse.
@@
$\eta(t)=\begin{cases}\frac{t}{w}\,\eta_{\max},&0\leq t<w\quad\text{(warm‑up)},\\[4.0pt] \eta_{\max},&w\leq t<p\quad\text{(plateau)},\\[2.0pt] \eta_{\max}\!\Bigl(1-\tfrac{t-p}{d}\Bigr),&p\leq t<p+d\quad\text{(cool‑down)},\end{cases}$
where $w$ denotes the warm‑up interval, pwp-w is the constant‑LR plateau, and $d$ represents the cool-down segment. This stable phase allows us to efficiently manage experiments by saving intermediate checkpoints and then running additional cool-down steps from those points, according to each budget. For the warmup, we allocate $5\,\%$ of the total training steps of the smallest budget (2EB), and we set the cool‑down steps to $20\,\%$ of the total training steps for each corresponding budget.
@@
Table 9 reports NLL on FineWeb-Edu validation set and few‑shot accuracy on six benchmarks. Our findings reveal a clear trend: more compute consistently leads to better models, evidenced by lower NLL and improved accuracy with higher FLOPs. However, weight-sharing alone in recursive models degraded performance compared to Vanilla, a clear trade-off for the reduced parameters. Crucially, token-routed MoR models overcome this shortcoming, catching up to and then surpassing Vanilla models from 360M parameters upward, all while utilizing only one-third of the parameters. This performance advantage persists at 730M and 1.7B parameter scales.
Table 9: Detailed results of isoFLOP analysis across three compute budgets. We evaluate negative log‑likelihood (NLL) on the FineWeb‑Edu validation set and few‑shot accuracy on six downstream tasks for four base model sizes (135M, 360M, 730M, 1.7B). Each model was initially trained up to 16.5EB and sliced back to 5EB and 2EB via mid‑training checkpoints using a trapezoid learning‑rate schedule. All models used three recursion steps. For MoR models, we use expert-choice routing and recursion-wise caching mechanisms. We highlight the best-performing model in each setting in gray.
@@
Appendix E Details of Experimental Settings for Throughput Measurement
@@
Appendix F Expanded Results of Parameter Sharing Strategy
This section complements the ablation in §4.1 by providing the full quantitative panorama behind 4(b). We revisit the four weight‑tying schemes—Cycle, Sequence, Middle‑Cycle, and Middle‑Sequence—on two base model scales (135M and 360M non‑embedding parameters) and two different recursion depths ($N_{r}=2$ and 3). All models were trained from scratch for 10B tokens under identical optimization hyperparameters. Validation NLL on FineWeb‑Edu and averaged few‑shot accuracy over six benchmarks are summarized in Table 10.
Middle‑Cycle is consistently the safest choice.
@@ Appendix F Expanded Results of Parameter Sharing Strategy
Figure 6: Validation negative log‑likelihood (lower is better) on FineWeb‑Edu for four parameter‑sharing strategies. Bars are grouped by model capacity (135M vs. 360M parameters). Middle‑Cycle consistently attains the lowest NLL, with its margin widening as either model size or depth increases. The horizontal dashed lines mark the untied (non-sharing) baselines: the lower red line represents the full capacity model, while the upper black line represents a parameter-matched reduced model with a footprint equal to the unique trainable parameter sizes of the recursive model.
[…]
Table 11: Comparison of Middle-Cycle and Cycle sharing strategies under the same number of unique parameters. All models use 730M scales (non-embedding parameters) as the base model with 3 recursion steps. The models are trained from scratch on 10B tokens using the Fineweb-Edu dataset.
Recursion Pretrain NLL $\downarrow$ Few-shot Accuracy $\uparrow$
Base Model Share Loop N-Emb $N_{L}$ $N_{tok}$ Fineweb LD HS PQ WG ARC MMLU Avg
Vanilla 730M Cycle 3 252M 10 10B 2.7573 28.72 37.31 65.56 52.01 39.94 27.85 41.90
Vanilla 730M M-Cycle 3 252M 1+8+1 10B 2.7552 29.75 37.73 66.00 52.49 40.49 27.50 42.32
Behavior under continued pre‑training (up‑training).
Table 12 extends the study by “up‑training” models—continuing from open-sourced SmolLM (Allal et al., 2024) checkpoints for an additional 5B tokens. Both Middle strategies demonstrate superior performance across all settings, and notably, they significantly outperform the reduced baseline models that are initialized in the same manner but without recursion. The other strategies reach a performance plateau earlier, suggesting that they have limited room for further improvement in capacity.
@@
Appendix G Expanded Results of Design Choices for Router
G.1 Details of Design Configurations
@@
G.2 Router Performance Evaluation Metrics
@@
[…]
@@
G.3 Extended Evaluation Results of Router Designs
@@
Appendix H Expanded Results of KV Cache Sharing Mechanism
H.1 Key Value Representation Trends in Recursive Transformers
Sharing KV caches across model depths has emerged as a promising approach to improve inference throughput in Vanilla Transformers (Brandon et al., 2024). This technique can reduce the memory footprint required for KV caches, enabling larger inference batch sizes. Significant speedups can be also achieved by skipping the KV projection and even prefill operations at shared depths, especially with Cycle strategy (Sun et al., 2024). Due to the high degree of freedom in Vanilla models—where trainable parameters can be well optimized for shared caches—these models exhibit only marginal performance drops when KV caches are shared between adjacent layers. In contrast, Recursive Transformers have far fewer parameters available for being optimized to tied KV states. Nevertheless, we hypothesize that similar patterns may emerge between shared blocks. To investigate this, we decomposed the KV states from pretrained Recursive Transformers into magnitude and directional components.
@@ H.1 Key Value Representation Trends in Recursive Transformers
H.2 Performance Comparison of KV Sharing Strategy
@@
[…]
@@
Appendix I Expanded Qualitative Results
I.1 Analysis on Adaptive Computation Paths
@@
Table 17: Visualization of the recursion depth for each subword token, with colors representing the number of recursion steps: 1, 2, and 3. This depth specifically indicates how many times recursion is applied at that token’s position to predict the subsequent token. Each row corresponds to a single sample, offering a clear illustration of the token-level recursion distribution in practice. We use an MoR model with $N_{r}=3$, auxiliary loss, and recursion-wise KV caching. This model is built on a 360M parameter base and trained on 30B tokens.
@@
I.2 Analysis on Router Weights